# Where to find Grokking in LLM Pretraining? Monitor Memorization-to-Generalization without Test

### v1

@@ Abstract
Our study, for the first time, verifies that grokking still happens in the pretraining of practical, large-scale foundation models, though different data may enter their grokking stages asynchronously. We further demystify grokking’s “emergence of generalization” by investigating the dynamics of LLMs’ internal states. Specifically, we find that training samples’ pathways (i.e., expert choices across layers) evolve from random, instance-specific to more structured and shareable between samples during grokking. Also, the complexity of a single sample’s pathway reduces despite the converged loss. These indicate a memorization-to-generalization conversion, providing a mechanistic explanation of the delayed generalization. In the study, we develop two novel metrics to quantify the distance between pathways and the complexity of a single pathway. We demonstrate their capabilities to predict the generalization improvement on diverse downstream tasks. They are efficient, simple to compute and solely dependent on training data. Hence, they have crucial practical value to model pretraining, enabling us to monitor the generalization performance without finetuning and test. Theoretically, we ground our findings in a spectrum analysis of the tangent kernel for one-layer MoE, showing that more structured pathways reduce model complexity and improve the generalization bound.
@@ 2 Related Work
Several works have focused on revealing the mechanisms behind grokking behavior. Liu et al. [11] demonstrated in a simplified setting that grokking results from structured representations emerging during training. Nanda et al. [16] further revealed that grokking can be attributed to the gradual amplification of structured mechanisms encoded within model weights. Using a single-layer ReLU network, Merrill et al. [14] associated grokking with subnetwork sparsity. Meanwhile, Humayun et al. [9] observed that decision boundaries become smoother during the grokking phase in CNN classification tasks. Beyond interpretability-focused research, Notsawo Jr et al. [17] identified oscillatory patterns in early training loss as predictors of subsequent grokking phenomena. Distinct from these works, our research provides an explanation of grokking within significantly larger LLMs through an interpretative analysis of MoE architectures. Additionally, we introduce practical indicators for monitoring generative behavior throughout LLM pretraining.
@@ 3.1 Domain-Level Analysis: Local Grokking Emerges Asynchronously across Domains
Figure 1: Converged groups of pretraining samples (accumulated) over training steps and test accuracy on downstream tasks across four domains. It shows the asynchronous memorization of local data groups and how it affects the global generalization: the test accuracy starts lately to rise sharply with the increasing converged training samples. But the starting time of grokking the induced gain vary across domains, explaining the difficulty to allocate global grokking. Math and coding tasks require more knowledge accumulation to generalize than commonsense and domain specific reasoning.
@@ 3.2 Group-Level Analysis: Later Grokking & Longer Delay for More Difficult Data
Figure 2: Training loss and test accuracy of four data groups over pretraining steps with local grokking. The pretraining data and test samples in each group are matched based on their semantic similarities. We observe local grokking, i.e., delayed rise of test accuracy after training loss plateaued, on every group, though their delayed steps and grokking period lengths vary. From left to right, later converged and memorized (harder) samples exhibit systematically longer delays.
@@ 4.3 Pathway Complexity Metrics as Monitors of Generalization
We define the global grokking point as the first checkpoint after which test accuracy exhibits a sustained upward trend (Figure 1), marking the onset of meaningful generalization. Our correlation analysis focuses on checkpoints after this grokking point—excluding the initial random-accuracy phase—to isolate the mechanisms driving generalization.
@@ 4.4 Theoretical Connections & Explanations
For a detailed theoretical setup, assumptions, and proof of Theorem A.1, please refer to Appendix A.
@@ A.2 Routing Kernel Definition
The structure of $\Theta_{\text{route}}(x,x^{\prime})$ in this fixed-routing setup provides crucial insights into how routing influences generalization. It is composed of two primary factors: the gating weight product $g_{j}(x)g_{j}(x^{\prime})$, and the individual expert NTKs $K_{f_{j}}(x,x^{\prime})$. The term $g_{j}(x)g_{j}(x^{\prime})$ signifies the importance of the $j$-th expert for both inputs $x$ and $x^{\prime}$, highlighting that two inputs are considered "similar" by the kernel if they are significantly routed to the same experts. If inputs $x$ and $x^{\prime}$ are routed to distinct sets of experts, their contribution to the sum for a given $j$ will be diminished. Simultaneously, $K_{f_{j}}(x,x^{\prime})$ captures the functional similarity learned by the $j$-th expert for inputs $x$ and $x^{\prime}$. Thus, the overall routing kernel measures a compounded similarity: inputs are similar not only if they are guided through similar pathways (weighted by $g_{j}(x)g_{j}(x^{\prime})$), but also if the specific experts they pass through exhibit similar functional behavior for those inputs ($K_{f_{j}}(x,x^{\prime})$). This decomposition underscores how fixed routing patterns (reflected in $g_{j}(x)$) set the stage for expert specialization and ultimately shape the global generalization properties of the MoE model.

### v2

@@ Abstract
Our study, for the first time, verifies that grokking still happens in the pretraining of practical, large-scale foundation models, though different data may enter their grokking stages asynchronously. We further demystify grokking’s “emergence of generalization” by investigating the dynamics of LLMs’ internal states. Specifically, we find that training samples’ pathways (i.e., expert choices across layers) evolve from random, instance-specific to more structured and shareable between samples. Also, the complexity of a single sample’s pathway reduces despite the converged loss. These indicate a memorization-to-generalization “knowledge digestion”, providing a mechanistic explanation of the delayed generalization. In the study, we develop two novel metrics to quantify the distance between pathways and the complexity of a single pathway. We demonstrate their capabilities to predict the generalization improvement on diverse downstream tasks. They are efficient, simple to compute and solely dependent on training data. Hence, they have crucial practical value to model pretraining, enabling us to monitor the generalization performance without finetuning and test. Theoretically, we ground our findings in a spectrum analysis of the tangent kernel for one-layer MoE, showing that more structured pathways reduce model complexity and improve the generalization bound.
@@ 2 Related Work
Several works have focused on revealing the mechanisms behind grokking behavior. Liu et al. [11] demonstrated in a simplified setting that grokking results from structured representations emerging during training. Nanda et al. [16] further revealed that grokking can be attributed to the gradual amplification of structured mechanisms encoded within model weights. Wang et al. [24] showed that even small Transformers trained from scratch on synthetic reasoning tasks exhibit a grokking-driven emergence of implicit reasoning, where generalization arises only after extended overfitting and internal structure formation. Using a single-layer ReLU network, Merrill et al. [14] associated grokking with subnetwork sparsity. Meanwhile, Humayun et al. [9] observed that decision boundaries become smoother during the grokking phase in CNN classification tasks. Beyond interpretability-focused research, Notsawo Jr et al. [17] identified oscillatory patterns in early training loss as predictors of subsequent grokking phenomena. Distinct from these works, our research provides an explanation of grokking within significantly larger LLMs through an interpretative analysis of MoE architectures. Additionally, we introduce practical indicators for monitoring generative behavior throughout LLM pretraining.
@@ 3.1 Domain-Level Analysis: Local Grokking Emerges Asynchronously across Domains
Figure 1: Converged groups of pretraining samples (accumulated) over training steps and test accuracy on downstream tasks across four domains. It shows the asynchronous memorization of data from four local domains and how it affects the corresponding generalization: each domain’s test accuracy starts to rise sharply after sufficient training samples converge. But the starting time of grokking and the induced gain vary across domains, resulting in domain-level local grokking rather than global grokking. Math and coding tasks require more knowledge accumulation to generalize than commonsense and domain-specific reasoning.
@@ 3.2 Group-Level Analysis: Later Grokking & Longer Delay for More Difficult Data
Figure 2: Training loss and test accuracy over pretraining steps of four data groups in the math domain with local grokking. The pretraining data and test samples in each group are selected and matched based on their semantic similarities. We observe local grokking on every group, i.e., delayed rise of test accuracy after training loss plateaued, though their delayed steps and grokking period lengths differ. From left to right, later converged and memorized (harder) samples exhibit systematically longer delays.
@@ 4.3 Pathway Complexity Metrics as Monitors of Generalization
We define the domain-level grokking point as the first checkpoint after which test accuracy for each domain exhibits a sustained upward trend (Figure 1), marking the onset of meaningful generalization. Our correlation analysis focuses on checkpoints after this grokking point—excluding the initial random-accuracy phase—to isolate the mechanisms driving generalization.
@@ 4.4 Theoretical Connections & Explanations
For a detailed theoretical setup, assumptions, and proof of Theorem 4.1, please refer to Appendix A.
@@ A.2 Routing Kernel Definition
The structure of $\Theta_{\text{route}}(x,x^{\prime})$ in this fixed-routing setup provides crucial insights into how routing influences generalization. It is composed of two primary factors: the gating weight product $g_{j}(x)g_{j}(x^{\prime})$, and the individual expert NTKs $K_{f_{j}}(x,x^{\prime})$. The term $g_{j}(x)g_{j}(x^{\prime})$ signifies the importance of the $j$-th expert for both inputs $x$ and $x^{\prime}$, highlighting that two inputs are considered "similar" by the kernel if they are significantly routed to the same experts. If inputs $x$ and $x^{\prime}$ are routed to distinct sets of experts, their contribution to the sum for a given $j$ will be diminished. Simultaneously, $K_{f_{j}}(x,x^{\prime})$ captures the functional similarity learned by the $j$-th expert for inputs $x$ and $x^{\prime}$. Thus, the overall routing kernel measures a compounded similarity: inputs are similar not only if they are guided through similar pathways (weighted by $g_{j}(x)g_{j}(x^{\prime})$), but also if the specific experts they pass through exhibit similar functional behavior for those inputs ($K_{f_{j}}(x,x^{\prime})$). This decomposition underscores how fixed routing patterns (reflected in $g_{j}(x)$) set the stage for expert specialization and ultimately shape the generalization properties of the MoE model.